a gas cylinder contains 320gms oxygen gas at 24.6atm pressure and 27celsius .what mass of oxygen gas would escape if first the cyclinder were heated to133 celsius and then the valve were held open until the gas pressure was 1.00atm,the temperature being maintained at 133 celsius??????????


 

Grade:11

2 Answers

gaurav gupta
12 years ago

when it was heated to 133celsius pressure and volume were constt.

so, nt=constt

(320/32)*300=(w/32)406

w=0.00423

when it was held open temp. and volume don't  change

so , p/n=constt

24.6/(0.00423/32)=1/(w1/32)

w1=0.000172

mass escaped=320-0.000172=319.9gms

don augusty
12 years ago

answer bahaut hi aasan hain  pv= nrt  will give u the solution for the answer